OpenStudy (anonymous):

David has a bag that holds 18 pens. 5 of the pens are ball point, 4 are gel pens, and 9 are felt tip marker pens. He reaches into the bag to pull out one pen. If every pen has an equal chance of being pulled out of David’s bag, what is the probability that he will pull out a gel pen?

OpenStudy (diyadiya):

4/18=2/9

OpenStudy (anonymous):

4/18=2/9 reduced

OpenStudy (anonymous):

4/19=2/9

OpenStudy (anonymous):

2/9

OpenStudy (anonymous):

David has a bag that holds 18 pens. 5 of the pens are ball point, 4 are gel pens, and 9 are felt tip marker pens. He reaches into the bag to pull out one pen. If every pen has an equal chance of being pulled out of David’s bag, what is the probability that he will pull out a gel pen?